Session: Feb 2012 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (Kaz Borkowski at Wed Mar 14 06:37:12 2012) Success:
No known problems.

Yebes (Daniel Valero at Wed Mar 14 06:49:09 2012) Success:
No known problems.

Hartebeesthoek (Jonathan Quick at Wed Mar 14 06:50:15 2012) Success:
No known problems, apart from the 3 hour wait for the last scan.

Medicina (Andrea Orlati at Wed Mar 14 06:57:05 2012) Success:
no known problems

Svetloe (T.Andreeva at Wed Mar 14 07:44:53 2012) Success:
Scan No0005 failed due to antenna problem

Urumqi (Wenjun Yang at Wed Mar 14 11:18:24 2012) Success:
"It is little snow.

Westerbork (Geert Kuper at Thu Mar 15 08:44:07 2012) Success:
No known problems.

Effelsberg (Uwe Bach at Thu Mar 15 10:21:43 2012) Minor failures: ANT
From 18:40 to 20:20 UT several breaks due to Azimuth control failures.

Onsala (Michael Lindqvist at Fri Mar 16 14:33:07 2012) Minor failures: RECO
Lost some time due to problem with a module.

Effelsberg (Uwe Bach at Mon Mar 19 08:34:35 2012) Success:
The last JIVE disk for this session had a wrong electronic VSN. The log quotes "MAY-0004" but the labelled VSN on the disk pack is "HAY-0004".

Jodrell_Bank (Alastair Gunn at Mon Mar 19 15:24:24 2012) Success:
No known problems.